What is the answer to 6t+5(5t-3)?
Found 2 solutions by mananth, alesvelasco:
Answer by mananth(16949)

6t+5(5t-3)
6t+25t-15
31t-15

Answer by alesvelasco(1)

6t+5(5t-3)=0
6t+25t-15=0
31t-15=0
31t=15
t=15/31